Thanks.. I think I did the wrong answer.
Was.. Determine the sum of the sussesion
An={3, 3/4, 3/16, 3/64...} (3/4., 3/16... are fracctions...)

Answer by fractalier(6550)

The sum of this infinite series is found by
S = a1 / (1-r)
S = 3 / (1 - (1/4))
S = 3 / (3/4) = 4